OpenStudy (anonymous):

Solve the system by the method of your choice. Identify systems with no solution and systems with infinitely many solutions, using set notation to express their solution sets. 4x - 3y = 6 -12x + 9y = -24

OpenStudy (anonymous):

Please explain how to solve so I can understand

OpenStudy (precal):

ok I would use the elimination method (one method you can apply to all systems)

OpenStudy (precal):

your goal is to eliminate x or y, your choice

OpenStudy (precal):

4x and -12x if you try to eliminate the x's you will need the same number I will choose 12 because 4 times 3 is 12

OpenStudy (precal):

I can create 12x by multiplying 3 to 4x

OpenStudy (precal):

that is one choice, if I choose to eliminate the y's

OpenStudy (precal):

what number would I use?

OpenStudy (anonymous):

-3y and 9y ...you could use 9

OpenStudy (aum):

Try multiplying the first equation by -3. What do you get?

OpenStudy (aum):

4x - 3y = 6 -12x + 9y = -24 Multiply first equation by -3: -12x + 9y = -18 Compare this to the second equation: The left hand sides are identical. But the right hand side is not. This is impossible. There are no values of x and y that can make -12x + 9y be -18 AND -24 at the same time. Therefore, this system of equations has no solution.

OpenStudy (anonymous):

what did you mean by the left hand side being identical?

OpenStudy (aum):

After multiply first equation by -3 I got: -12x + 9y = -18 Compare this to the second equation: -12x + 9y = -24 The left hand side in both equations is the same: -12x + 9y How can the same thing (-12x+9y) be two different things: -18 AND -24? So no solutions because the equations are incompatible.

OpenStudy (anonymous):

Understood...thank you very much. ♥